# Photon propagation in a charged Bose-Einstein condensate

### v1

@@ Abstract
We consider the propagation of photons in the background of a Bose-Einstein (BE) condensate of a charged scalar field, by extending a method recently proposed to treat the propagation of fermions in a BE condensate. We determine the dispersion relations of the collective modes of the system, as well as the photon polarization tensor and the dielectric constant that result after the symmetry breaking associated with the BE condensation in the model. Two modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. The other two modes, which we denote as the (±)plus-or-minus($\pm$)( ± ) modes, are combinations of the longitudinal photon and the massive scalar field. The dispersion relation of the ()(-)( - ) mode decreases as a function of the momentum in a given range, and the corresponding group velocity is negative in that range. We also determine the wavefunctions of the (±)plus-or-minus($\pm$)( ± ) modes, which can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system. The results can be useful in various physical contexts that have been considered in the literature involving the electrodynamics of a charged scalar BE condensate.
@@ 1 Introduction and motivation
Our motivation in the present work is to consider, in a similar vein, the propagation of photons in a thermal background that contains a BE condensate of a charged scalar. The problem of the propagation of photons in a medium is a classic subject which in the language of plasma physics can be treated within the formalism of the kinetic Boltzmann equation, and in other contexts and/or more generally using the framework of thermal field theory (TFT). Various aspects and applications of the problem of photon propagation in a BE condensate of the charged scalars have been studied in previous recent works, in various contexts and using different approaches, for example the electrodynamics of fermions and charged scalar bosons[10], Cherenkov effect[11], exciton-polaritons[12], the BE condensation of photons[13, 14] and the reduction of the photon velocity in atomic gases[15].
@@
Our goal in the present work is to apply the same technique used in Ref. [2], to the case in which the scalar is electromagnetically charged and study the propagation of photons in the BE condensate background of such scalars. The field theoretical method we used in Ref. [2] to treat the BE condensate, and apply here, has been discussed by various authors[16, 17, 18]. As we will see, the implementation proposed here to treat the propagation of photons allows us to determine succinctly the dispersion relations of the collective modes that result after the symmetry breaking associated with the BE condensation in the model. We are also able to derive the polarization tensor and the dielectric constant. The method also allows us to determine in a very efficient way the mixing parameters between the longitudinal photon (plasmon) and the physical scalar field that make up the propagating modes involving the plasmon. These are relevant for the calculation of the thermal corrections to the dispersion relations (e.g., imaginary parts due to damping effects) and the other physical quantities due to interactions with the thermal excitations of the BE condensate. The results described above are obtained in a general way, within the context of the model, but not tied to any specific application. In this way they can be useful and pave the way for further development in various contexts and applications such as those considered in the references mentioned above.
@@
The plan of the presentation describing the present work is as follows. In Section 2 we summarize the essential features and the symmetry breaking mechanism of the scalar BE condensation model discussed in Ref. [2], which serves as the basis and motivation for the extension we present in Section 3 to the charged scalar case. In Section 3 we discuss various qualitative properties of the model that are useful for the interpretation, such as current conservation, symmetry breaking and the spectrum. The propagating modes and the corresponding dispersion relations are determined in Section 4. As we show, two modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. The other two modes, which we denote as the (±)plus-or-minus($\pm$)( ± ) modes, are combinations of the longitudinal photon and the massive scalar field. Various distinguishing aspects of the results obtained are discussed in Section 5. A particularly noticeable one is the fact that the dispersion relation of the ()(-)( - ) mode decreases as a function of the momentum in a given range, and the corresponding group velocity is negative in that range. In Section 6 we obtain the expression for the photon polarization tensor, or self-energy. This provides an alternative way to obtain the dispersion relations, but more importantly it has also the virtue that it provides an expression for the dielectric function of the system, which is useful on its own merits, for example for studying also the response of the system to external fields. Finally in Section 7 we determine the mixing parameters of the spinors of the (±)plus-or-minus($\pm$)( ± ) modes. These parameters can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system. In Section 8 we summarize our conclusions and outlook.
@@ 2 Scalar BE condensate model
As a preamble to the presentation of the model for the charged scalar coupled to the photon, we first summarize the essential features of the scalar BE condensation model discussed in Ref. [2], as well as its justification.
@@
Expanding the $d$ term in Eq. (2.1), the Lagrangian is
@@
Substituting Eqs. (2.9) and (2.12) in Eq. (2.1) we obtain the Lagrangian for the scalar excitations $\phi_{1,2}$,
@@
The Lagrangian, in momentum space, can then be written in the form
@@
$D=(\omega^{2}-\kappa^{2})(\omega^{2}-\kappa^{2}-m^{2}_{1})-4\mu^{2}\omega^{2}\,,$ (2.22)
where we have used Eq. (2.5). The dispersion relations obtained in this way can be written in the form
$\omega^{2}_{\pm}(\kappa)=\kappa^{2}+\frac{1}{2}m^{2}_{+}\pm\sqrt{\frac{1}{4}m^{4}_{+}+4\mu^{2}\kappa^{2}}\,,$ (2.23)
@@
$m^{2}_{+}=m^{2}_{1}+4\mu^{2}\,.$ (2.24)
The zero mass mode, i.e., $\omega_{-}(0)=0$, is the realization of the Goldstone mode associated with the breaking of the global $U(1)$ symmetry.
As discussed in Ref. [2], the motivation for considering this model is based on the following result: the calculation of the effective potential $V_{\text{eff}}(T,\mu)$ for $\phi$ can be carried out in TFT using $\mu=0$ in the partition (and/or distribution) function, but using the $\mu$-dependent Lagrangian $L_{\phi}$ defined in Eq. (2.1). Then neglecting the $T$-dependent terms (that is, at zero temperature), $V_{\text{eff}}(0,\mu)$ is simply the $U$ potential given in Eq. (2.7). In other words, calculations with the $\phi$ propagators, $\phi$ Lagrangian and partition function for $\phi$,
@@
give the same result as calculating them with the Lagrangian for $\phi^{\prime}$ obtained by using $\partial_{\mu}\rightarrow D_{\mu}=\partial_{\mu}-i\mu u_{\mu}$ and using
@@ 3.1 Lagrangian
We extend the model by assuming that the $\phi$ field is electrically charged and interacts with electromagnetic field in the usual way. The Lagrangian is then
@@
3.2 Current
It will be useful for later purposes to identify the electromagnetic current, in particular its zeroth component or the charge density. The electromagnetic current, defined as
$j^{(em)}_{\mu}=-\frac{\partial L_{\phi A}}{\partial A^{\mu}}\,,$ (3.6)
is then given by
$j^{(em)}_{\mu}=iqJ_{\mu}\,,$ (3.7)
where
$J_{\mu}=\phi^{\ast}\hat{D}_{\mu}\phi-(\hat{D}_{\mu}\phi)^{\ast}\phi\,.$ (3.8)
For consistency we verify that it is conserved as a consequence of the equation of motion.
For the equation of motion we first get
$\frac{\partial L_{\phi A}}{\partial\phi^{\ast}}=i\omega^{\mu}(\hat{D}_{\mu}\phi)-\frac{\partial V_{\phi}}{\partial\phi^{\ast}}\,,$ (3.9)
and
$\frac{\partial L_{\phi A}}{\partial(\partial^{\mu}\phi^{\ast})}=\hat{D}_{\mu}\phi\,.$ (3.10)
The Lagrange equation
$\partial^{\mu}\frac{\partial L_{\phi A}}{\partial(\partial^{\mu}\phi^{\ast})}-\frac{\partial L_{\phi A}}{\partial\phi^{\ast}}=0\,,$ (3.11)
then gives the equation of motion
$\partial^{\mu}(\hat{D}_{\mu}\phi)-i\omega^{\mu}(\hat{D}_{\mu}\phi)+\frac{\partial V_{\phi}}{\partial\phi^{\ast}}=0\,,$ (3.12)
or
$\hat{D}^{2}\phi+\frac{\partial V_{\phi}}{\partial\phi^{\ast}}=0\,.$ (3.13)
To verify
$\partial\cdot J=0\,,$ (3.14)
from Eq. (3.8) we get
$\partial\cdot J=(\partial\phi)^{\ast}\cdot\hat{D}\phi+\phi^{\ast}\partial\cdot\hat{D}\phi-(\partial\cdot\hat{D}\phi)^{\ast}\phi-(\hat{D}\phi)^{\ast}\cdot\partial\phi\,.$ (3.15)
Putting
$\partial_{\mu}=\hat{D}_{\mu}+i\omega_{\mu}\,,$ (3.16)
the four terms with the $i\omega_{\mu}$ cancel in pairs, the two terms of the form $(\hat{D}\phi)^{\ast}(\hat{D}\phi)$ cancel as well, and we are left with
$\partial\cdot J=\phi^{\ast}\hat{D}^{2}\phi-(\hat{D}^{2}\phi)^{\ast}\phi\,,$ (3.17)
which is zero by the equation of motion.
@@ 3.3 Spectrum of the model
The physical picture that emerges is this: the field $\theta$ becomes the longitudinal component of $V$, and we end up with two fields, $V$ and $\rho$. The longitudinal component of $V$ is mixed with the scalar field $\rho$ (the term $\rho v\cdot V)$. What remains is to find the dispersion relations of the propagating modes, which appear to be similar that the dispersion relations of the ϕ±subscriptitalic-$\phi_{\pm}$ of the BE condensate model discussed in Section 2. As we show the consequence is that the propagating modes with definite dispersion relations involve a superposition of the longitudinal component of $V$ and the $\rho$.
The following comment is in order before proceeding. In Eq. (3.20) there is a term linear in $V_{\mu}$ of the form $2q\phi_{0}\rho v\cdot V=(2q\mu\phi^{2}_{0})u\cdot V$. This term is a tadpole, meaning that the ground state has a net contribution to the charge due to $\phi_{0}$. Indeed from the Eq. (3.7), it follows that the symmetry breaking mechanism produces a net charge density given by
$j^{(em)0}=qn_{0}\,,$ (3.26)
with
$n_{0}=2\mu\phi^{2}_{0}\,.$ (3.27)
Physically, the total charge of the system must be zero, which is assured by assuming that the $\phi$ background is superimposed on an oppositely charged background to render the total system electrically neutral. Thus for practical purposes we can disregard that term.
@@ 4.1 The evolution equations
If it were not for the last term, this is the usual free Lagrangian for a vector with squared mass $m^{2}_{V}$ and a scalar with squared mass $m^{2}_{1}$. However the last term mixes the scalar with the longitudinal components of $V^{\mu}$.
@@
$\tilde{u}_{\mu}=\tilde{g}_{\mu\nu}u^{\nu}=u_{\mu}-\frac{(k\cdot u)k_{\mu}}{k^{2}}\,.$ (4.6)
@@
Thus the quadratic part of the Lagrangian is
@@
which is the same that is obtained from Eq. (4.12). The picture that emerges is that the transverse modes develop a standard degenerate dispersion relation for the two polarizations, while the longitudinal mode $V_{L}$ mixes with the scalar $\rho$. Our next task is to solve the evolution equations we have obtained, to find the dispersion relations and the corresponding mixing parameters $V_{L}$ and $\rho$ eigenmodes.
@@ 4.2 Dispersion relations
At $\kappa=0$ (remembering that $\tilde{u}^{2}=-\kappa^{2}/k^{2}$), the equations in Eq. (4.1) decouple,
@@
$(\omega^{2}-m^{2}_{1}-4\mu^{2})\rho$ $=$ $0\,,$ (4.15)
so we can identify the plasma frequency as
$\omega^{2}_{p0}=m^{2}_{V}=q^{2}\phi^{2}_{0}=\frac{q^{2}(\mu^{2}-m^{2})}{\lambda}\,.$ (4.16)
The $\rho$ mode has an effective mass
$\omega^{2}_{\rho 0}=m^{2}_{1}+4\mu^{2}=m^{2}_{+}\,,$ (4.17)
which coincides with the mass of the “$+$” mode in the un-gauged model (see Eq. (2.24)).
@@
or
$(k^{2}-m^{2}_{V})(k^{2}-m^{2}_{1})-\frac{4\mu^{2}}{k^{2}}\left[(k\cdot u)^{2}(k^{2}-m^{2}_{V})+m^{2}_{V}\kappa^{2}\right]=0\,.$ (4.19)
We now use the definition of $\kappa$,
$(u\cdot k)^{2}=k^{2}+\kappa^{2}\,,$ (4.20)
to rewrite the term in square brackets,
$(k\cdot u)^{2}(k^{2}-m^{2}_{V})+m^{2}_{V}\kappa^{2}=k^{2}(k^{2}-m^{2}_{V}+\kappa^{2})\,,$ (4.21)
so that the equation becomes
$(k^{2}-m^{2}_{V})(k^{2}-m^{2}_{1})-4\mu^{2}\left[k^{2}-m^{2}_{V}+\kappa^{2}\right]=0\,,$ (4.22)
which write it as
$(k^{2})^{2}-(m^{2}_{V}+m^{2}_{+})k^{2}+m^{2}_{V}m^{2}_{+}-4\mu^{2}\kappa^{2}=0\,,$ (4.23)
where $m^{2}_{+}$ is defined in Eq. (2.24) (which coincides with the squared mass of the “$+$” mode in the un-gauged model and with effective squared mass of the $\rho$ as indicated in Eq. (4.17)). This is a quadratic equation for $k^{2}$, with solutions
$\omega^{2}=\kappa^{2}+\frac{1}{2}(m^{2}_{V}+m^{2}_{+})\pm\left[\frac{1}{4}(m^{2}_{V}-m^{2}_{+})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,.$ (4.24)

### v2

@@ Abstract
We consider the propagation of photons in a model of a charged scalar Bose-Einstein (BE) condensate. We determine the dispersion relations of the collective modes, as well as the photon polarization tensor and the dielectric constant in the model. Two modes correspond to the transverse photon polarizations, with dispersion relations of the usual form for transverse photons in a plasma. The other two modes, denoted as the (±)plus-or-minus($\pm$)( ± ) modes, are combinations of the longitudinal photon and the massive scalar field. Their dispersion relations behave very differently as functions of momentum. The (+)(+)( + ) mode dispersion relation increases steadily and remains greater than the momentum as the momentum increases. The dispersion relation of the ()(-)( - ) mode decreases in a given momentum range, with the group velocity being negative in that range, while in another range it increases steadily but remains smaller than the momentum, akin to the situation in a medium with an index of refraction greater than 1. We consider the non-relativistic limit of the (±)plus-or-minus($\pm$)( ± ) dispersion relations and discuss some aspects of the results. We also determine the wavefunctions of the (±)plus-or-minus($\pm$)( ± ) modes, which are useful to obtain the corrections to the dispersion relations, e.g., imaginary parts due to the damping effects and/or the effects of scattering, due to the interactions with the excitations of the system. The results can be useful in various physical contexts that have been considered in the literature involving the electrodynamics of a charged scalar BE condensate.
@@ 1 Introduction and motivation
Our motivation in the present work is to consider, in a similar vein, the propagation of photons in a thermal background that contains a BE condensate of a charged scalar. The problem of the propagation of photons in a medium is a classic subject which in the language of plasma physics can be treated within the formalism of the kinetic Boltzmann equation, and in other contexts and/or more generally using the framework of thermal field theory (TFT). Various aspects and applications of the problem of photon propagation in a BE condensate of the charged scalars have been studied in previous recent works, in various contexts and using different approaches, for example the electrodynamics of fermions and charged scalar bosons at low temperature[10], heavy-ion collisions[11], exciton-polaritons[12], the BE condensation of photons[13, 14] and the Cerenkov effect[15], and the reduction of the photon velocity[16], in atomic gases.
@@
Our goal in the present work is to extend the technique used in Ref. [2] to the case in which the scalar is electromagnetically charged and study the propagation of photons in the BE condensate background of such scalars. The field theoretical method we used in Ref. [2] to treat the BE condensate, which we use here, has been discussed by various authors[17, 18, 19, 20]. As we will see, this method allows us to determine succinctly the dispersion relations of the propagating modes after the symmetry breaking associated with the BE condensation in the model, as well as the corresponding polarization tensor and dielectric constant. The method also allows us to determine in a very efficient way the mixing parameters between the longitudinal photon (plasmon) and the physical scalar field that make up the propagating modes involving the plasmon. These are relevant for the calculation of thermal corrections to the dispersion relations (e.g., imaginary parts due to damping effects) and the other physical quantities due to interactions with the thermal excitations of the BE condensate. The results described above are obtained in a general way, within the context of the model, but not tied to any specific application. In this way they can be useful and pave the way for further development in various contexts and applications such as those considered in the references mentioned above. While the principles and techniques used here to treat the charged BE condensate share some common ingredients with those of Ref. [20], our scope, discussion and outlook for applications are very different.
@@
The plan of the presentation describing the present work is as follows. In Section 2 we summarize the essential features and the symmetry breaking mechanism of the scalar BE condensation model discussed in Ref. [2], which serves as the basis and motivation for the extension we present in Section 3 to the charged scalar case. In Section 3 we discuss various qualitative properties of the model that are useful for the interpretation, such as current conservation, symmetry breaking and the spectrum. The propagating modes and the corresponding dispersion relations are determined in Section 4. As we show, two modes correspond to the transverse photon polarizations, and their dispersion relations have the usual form of the transverse photons in a plasma. The other two modes, which we denote as the (±)plus-or-minus($\pm$)( ± ) modes, are combinations of the longitudinal photon and the massive scalar field. Various distinguishing aspects of the results obtained are discussed in Section 5. While the dispersion relation of the (+)(+)( + ) mode increases steadily and remains greater than the momentum as the momentum increases, the dispersion relation of the ()(-)( - ) mode reveals some special features. A particularly noticeable one is the fact that it decreases in a given momentum range, and the corresponding group velocity is negative in that range. In another range, it increases steadily for values of the momentum above a given threshold, but such that it remains smaller than the momentum, reminiscent of a medium with index of refraction greater than 1. We also consider the non-relativistic limit of the dispersion relations and discuss some aspects of the results obtained. In particular, under some conditions, the dispersion relation of the ()(-)( - ) mode reduces to the expression for the dispersion relation that has been obtained in non-relativistic models of the BE condensation of charged scalars[21, 22]. In Section 6 we obtain the expression for the photon polarization tensor, or self-energy. On one hand, this provides an alternative way to obtain the dispersion relations. More importantly, it has the virtue that it gives the contribution to the dielectric function of the system due to the BE condensate, which is useful on its own merits, for example for studying also the response of the system to external fields. Finally in Section 7 we determine the mixing parameters of the spinors of the (±)plus-or-minus($\pm$)( ± ) modes. These parameters can be used to obtain the corrections to the dispersion relations (e.g., imaginary parts due the damping effects) and/or the effects of scattering, due to the interactions with the excitations of the system. In Section 8 we summarize our conclusions and outlook.
@@ 2 Scalar BE condensate model
As a preamble to the presentation of the model for the charged scalar coupled to the photon, we first summarize the essential features of the scalar BE condensation model as well as its justification.
@@
Expanding the $d$ term in Eq. (2.1), the Lagrangian density is
@@
Substituting Eqs. (2.9) and (2.12) in Eq. (2.1) we obtain the Lagrangian density for the scalar excitations $\phi_{1,2}$,
@@
The Lagrangian density, in momentum space, can then be written in the form
@@
$D=(\omega^{2}-\kappa^{2})(\omega^{2}-\kappa^{2}-m^{2}_{1})-4\mu^{2}\omega^{2}\,.$ (2.22)
In Eq. (2.22) we have used Eq. (2.5) and introduced the variables $\omega$ and $\kappa$,
$\omega$ $=$ $u\cdot k\,,$
$\kappa$ $=$ $\sqrt{\omega^{2}-k^{2}}\,.$ (2.23)
In the rest frame of the medium
$k^{\mu}=(\omega,\vec{\kappa})\,,$ (2.24)
with $\kappa=|\vec{\kappa}|$. The dispersion relations obtained in this way can be written in the form
$\omega^{2}_{\pm}(\kappa)=\kappa^{2}+\frac{1}{2}m^{2}_{\rho}\pm\sqrt{\frac{1}{4}m^{4}_{\rho}+4\mu^{2}\kappa^{2}}\,,$ (2.25)
@@
$m^{2}_{\rho}=m^{2}_{1}+4\mu^{2}\,,$ (2.26)
and they satisfy
$\omega_{+}(0)$ $=$ $m_{\rho}\,,$
$\omega_{-}(0)$ $=$ $0\,.$ (2.27)
The combinations of $\phi_{1,2}$ that have the definite dispersion relations $\omega_{\pm}$ can be obtained from Eq. (2.20). The mode corresponding to $\omega_{-}$ is the realization of the Goldstone mode associated with the breaking of the global $U(1)$ symmetry. However, its dispersion relation is not the usual $\omega=\kappa$, but the one given by $\omega_{-}(\kappa)$ in Eq. (2.25).
As discussed in Ref. [2], the motivation for considering this model is based on the following result: the calculation of the effective potential $V_{\text{eff}}(T,\mu)$ for $\phi$ can be carried out in TFT using $\mu=0$ in the partition (and/or distribution) function, but using the $\mu$-dependent Lagrangian density $L_{\phi}$ defined in Eq. (2.1). Then neglecting the $T$-dependent terms (that is, at zero temperature), $V_{\text{eff}}(0,\mu)$ is simply the $U$ potential given in Eq. (2.7). In other words, calculations with the $\phi$ propagators, $\phi$ Lagrangian density and partition function for $\phi$,
@@
give the same result as calculating them with the Lagrangian density for $\phi^{\prime}$ obtained by using $\partial_{\mu}\rightarrow D_{\mu}=\partial_{\mu}-i\mu u_{\mu}$ and using
@@ 3.1 Lagrangian
We extend the model by assuming that the $\phi$ field is electrically charged and interacts with electromagnetic field in the usual way. The Lagrangian density is then
@@ 3.2 Spectrum of the model
The physical picture that emerges is this: the field $\theta$ becomes the longitudinal component of $V$, and we end up with two fields, $V$ and $\rho$. The longitudinal component of $V$ is mixed with the scalar field $\rho$ (the term $\rho v\cdot V)$. As we show, the consequence is that the propagating modes with definite dispersion relations involve a superposition of the longitudinal component of $V$ and the $\rho$.
The following comment is in order before proceeding. In Eq. (3.8) there is a term linear in $V_{\mu}$ of the form $q\phi^{2}_{0}v\cdot V=(q\mu\phi^{2}_{0})u\cdot V$. This term is a tadpole, meaning that the ground state has a net contribution to the charge due to $\phi_{0}$. Indeed, the definition of the electromagnetic current,
$j^{(em)}_{\mu}=-\frac{\partial L_{\phi A}}{\partial A^{\mu}}\,,$ (3.14)
gives
$j^{(em)}_{\mu}=iqJ_{\mu}\,,$ (3.15)
where
$J_{\mu}=\phi^{\ast}\hat{D}_{\mu}\phi-(\hat{D}_{\mu}\phi)^{\ast}\phi\,,$ (3.16)
and it is straightforward to verify that $j^{(em)}_{\mu}$ is conserved, as it should be. It then follows that the symmetry breaking mechanism produces a net current density given by
$j^{(em)}_{\mu}=q\mu\phi^{2}_{0}u_{\mu}\,.$ (3.17)
Physically, the total charge of the system must be zero, which is assured by assuming that the $\phi$ background is superimposed on an oppositely charged background to render the total system electrically neutral. Thus for practical purposes we can disregard that tadpole term.
@@ 4.1 The evolution equations
If it were not for the last term, this is the usual free Lagrangian density for a vector with squared mass $m^{2}_{V}$ and a scalar with squared mass $m^{2}_{1}$. However the last term mixes the scalar with the longitudinal components of $V^{\mu}$.
@@
$\tilde{u}_{\mu}=\tilde{g}_{\mu\nu}u^{\nu}$ $=$ $u_{\mu}-\frac{(k\cdot u)k_{\mu}}{k^{2}}\,,$
$\sqrt{-\tilde{u}^{2}}$ $=$ $\frac{\kappa}{\sqrt{k^{2}}}\,.$ (4.6)
@@
Thus the quadratic part of the Lagrangian density is
@@
which is the same that is obtained from Eq. (4.12). The picture that emerges is that the transverse modes develop a standard degenerate dispersion relation for the two polarizations, while the longitudinal mode $V_{L}$ mixes with the scalar $\rho$. Our next task is to solve the evolution equations we have obtained, to find the dispersion relations and the corresponding mixing parameters $V_{L}$ and $\rho$ of the eigenmodes.
@@ 4.2 Dispersion relations
Using Eq. (4.1), it follows that at $\kappa=0$ the equations in Eq. (4.1) decouple,
@@
$(\omega^{2}-m^{2}_{\rho})\rho$ $=$ $0\,,$ (4.15)
where
$m^{2}_{\rho}=m^{2}_{1}+4\mu^{2}\,.$ (4.16)
Thus we can identify the plasma frequency as
$\omega^{2}_{p0}=m^{2}_{V}=q^{2}\phi^{2}_{0}=\frac{q^{2}(\mu^{2}-m^{2})}{\lambda}\,,$ (4.17)
while the $\rho$ mode has mass $m^{2}_{\rho}$, which coincides with the mass of the “$+$” mode in the un-gauged model (see Eq. (2.26)).
@@
which using Eq. (2) can be written as
$(k^{2}-m^{2}_{\rho})(k^{2}-m^{2}_{V})-4\mu^{2}\kappa^{2}=0\,.$ (4.19)
This is a quadratic equation for $k^{2}$, with solutions
$\omega^{2}=\kappa^{2}+\frac{1}{2}(m^{2}_{\rho}+m^{2}_{V})\pm\left[\frac{1}{4}(m^{2}_{\rho}-m^{2}_{V})^{2}+4\mu^{2}\kappa^{2}\right]^{\frac{1}{2}}\,.$ (4.20)